OpenStudy (anonymous):

write this ratio in simplest form 24 to 14 7 to 12 12 to 7 6 to 7 7 to 6

OpenStudy (anonymous):

abc or d

OpenStudy (lilg132):

12 to 7

OpenStudy (hoblos):

24/14 = 12/7

OpenStudy (anonymous):

\[\left\{\frac{12}{7},\frac{7}{12},\frac{12}{7},\frac{6}{7},\frac{7}{6}\right\} \]"12 to 7"